# Jinhong Li1    Qiuping Wang1    Shujie Han2    Patrick P. C. Lee1

### v2

@@ 5.1 Methodology
5.2 Microbenchmarks
@@
We examine the performance of ZapRAID in writes, normal reads, degraded reads, crash recovery, and full-drive recovery. In our default setup, we configure the logical space capacity as 200 GiB and set $G=256$ for ZapRAID in its group-based data layout. First, we focus on one open segment with various chunk sizes in Exp#1-Exp#6. We do not consider garbage collection (by limiting the total write size to be less than the logical space capacity) and keep the L2P table entirely in memory. We also consider the impact of various group sizes (Exp#3), the impact of RAID schemes (Exp#4), as well as the crash recovery and full-drive recovery performance with different logical space capacities (Exp#5). Furthermore, we consider multiple open segments in Exp#7-Exp#9, including the write performance under multiple open segments (Exp#7), the overhead of garbage collection (Exp#8), and the overhead of L2P table offloading (Exp#9).
Exp#1 (Write performance).
@@
We first evaluate the write performance of ZapRAID on a single open segment. We use Flexible IO Tester (FIO) (v3.35) [1] to generate a write-only workload that issues random writes of 64 GiB of data. We set the queue depth as 64 to saturate the system parallelism (Exp#6). We compare ZapRAID with RAIZN-SPDK, ZoneWrite-Only, and ZoneAppend-Only.
@@
First, we issue write requests of 4 KiB, 8 KiB, and 16 KiB, where the request size is the same as the chunk size. Figures 6(a)-6(c) show the throughput as well as the median and 95th-percentile latencies. Compared with ZoneWrite-Only, for 4-KiB and 8-KiB writes, ZapRAID achieves 72.8% and 77.2% higher write throughput, 43.8% and 44.5% lower median latency, and 36.4% and 34.1% lower 95th-percentile latency, respectively. This shows that ZapRAID improves the write performance by exploiting the intra-zone parallelism using Zone Append for small writes. On the other hand, for 16-KiB writes, ZapRAID has similar write throughput and median latency, but its 95th-percentile latency is 3.80×3.80$\times$3.80 × compared with that of ZoneWrite-Only. A possible reason is that Zone Append has high computational overhead in the ZNS SSD firmware, which incurs higher delays in processing individual Zone Append commands.
Note that in Figure 6(a), the throughput of 4-KiB writes under ZoneWrite-Only (910.8 MiB/s) is higher than that reported in our conference paper [71] (662.4 MiB/s). The reason is that we now optimize the implementation of ZoneWrite-Only by parallelizing the RAID operations (e.g., parity computations and stripe metadata updates) of the current stripe and the writes of the next stripe (as opposed to serially executing the operations). Now, the throughput is close to 3$\times$ the write throughput of Zone Write to an open zone (337.6 MiB/s in §2.1). This implies that the main bottleneck of ZoneWrite-Only lies in I/Os rather than computations.
@@
Compared with ZoneAppend-Only, ZapRAID achieves similar throughput and latencies for all write sizes, even though ZapRAID needs to wait for the completion of all Zone Append commands in a stripe group before issuing the writes for the next stripe group. Note that ZapRAID has much less memory usage (§3.2) and much lower degraded read latency (Exp#3) than ZoneAppend-Only.
@@
Compared with RAIZN-SPDK, all other three schemes show much higher performance on a single segment. The reasons are two-fold. First, RAIZN does not support Zone Append for intra-zone parallelism. Second, it uses partial parity updates to acknowledge individual write requests instead of issuing full-stripe writes as in the other three schemes. This limits the concurrency of write requests and hence the overall write performance. In Exp#7, we further compare ZapRAID with RAIZN-SPDK under multiple open segments.
Refer to caption Refer to caption Refer to caption
(a) Write throughput; (request size $=$ chunk size) (b) Median write latency; (request size $=$ chunk size) (c) 95p write latency; (request size $=$ chunk size)
Refer to caption Refer to caption Refer to caption
(d) Write throughput; (request size $=$ 4 KiB) (e) Median write latency; (request size $=$ 4 KiB) (f) 95p write latency; (request size $=$ 4 KiB)
@@
Figure 6: Exp#1 (Write performance on a single open segment). We show the results when the request size is equal to the chunk size (figures (a)-(c)) and when the request size is fixed as 4 KiB (figures (d)-(f)).
@@
We also issue write requests of a fixed size of 4 KiB, while varying the chunk size, to show the impact of the chunk size on write performance. Figures 6(d)-6(f) show the results. ZapRAID still maintains the high throughput and low median latencies for various chunk sizes, but its 95th-percentile latency for 16-KiB chunks is 50.7% higher than that for 8-KiB chunks. Similar results also appear in ZoneAppend-Only. The result indicates that Zone Append is mainly beneficial for small writes. It also justifies why we choose Zone Write over Zone Append in hybrid data management (§3.3).
Exp#2 (Normal and degraded read performance).
@@
We compare the performance of both normal reads and degraded reads in Log-RAID and ZapRAID. We vary the chunk size and set the read size to be the same as the chunk size. For each chunk size, we first fill up the 200-GiB logical address space to ensure that every logical block being read is indeed physically stored. We then use FIO to generate a read-only workload that issues random reads to 16 GiB of data using the randread option in FIO. We set the queue depth as one, so as to focus on the performance of individual read requests and exclude the interference among requests. We consider three cases: normal reads (NR), degraded reads under the static mapping in Log-RAID (DR-Log-RAID) (§2.3), and degraded reads under the group-based data layout in ZapRAID (DR-ZapRAID). Note that both Log-RAID and ZapRAID have the same workflow for normal reads. To evaluate degraded reads, we fail a drive and issue reads to the lost blocks of the failed drive.

### v3

@@ 5.1 Methodology
Our evaluation is driven by the following research questions:
  • How does ZapRAID perform on real applications in terms of writes, degraded reads, and memory usage for index management, when compared to ZoneWrite-Only and ZoneAppend-Only? (RQ1)
  • How does ZapRAID perform under realistic cloud storage traces that are composed of a mix of small and large writes? (RQ2)
  • How does ZapRAID perform across different aspects under microbenchmarks, including writes, normal reads, degraded reads, crash recovery, and full-drive recovery? (RQ3)
5.2 Real-Application Experiments (RQ1)
We start our evaluation with real applications to demonstrate the end-to-end performance benefits of ZapRAID, in comparison with ZoneWrite-Only and ZoneAppend-Only (similar benchmarks are also reported in [31]). We exclude RAIZN-SPDK, as it shows poor performance from our microbenchmarks (§5.4).
To host Linux applications, we export our ZapRAID prototype as an NVMe-over-Fabrics (NVMe-oF) target with TCP transport [64] and deploy the NVMe-oF host in the same testbed machine, so that the prototype serves as a Linux block device. We allocate the block device with a logical space capacity of 1 TiB, format the block device as an EXT4 file system, and run different types of applications atop the file system. Since the application write traffic may exceed the logical capacity, we reserve an additional 25% of the physical capacity for garbage collection.
We evaluate ZapRAID’s performance on both single-segment and multi-segment settings. For multi-segment settings, we vary the numbers of small-chunk and large-chunk segments (§3.3). We fix the small- and large-chunk sizes at $C_{s}=8$ KiB and $C_{l}=16$ KiB, respectively, and keep the total number of open segments as four (i.e., $N_{s}+N_{l}=4$). Here, we focus on four representative settings of open segments: (i) a single open segment with a chunk size of 4 KiB, (ii) a single open segment with a chunk size of 16 KiB, (iii) four open segments with $(N_{s},N_{l})=(2,2)$ (i.e., two small-chunk segments and two large-chunk segments), and (iv) four open segments with $(N_{s},N_{l})=(1,3)$ (i.e., one small-chunk segment and three large-chunk segments). We also present extensive evaluation results of various open-segment configurations in microbenchmarks in §5.4.
Workloads. We evaluate three types of application workloads: file writes, key-value workloads, and concurrent workloads.
File writes. We evaluate ZapRAID in file writes using Sysbench (v1.0.20) [66] to generate continuous write patterns. We first create 128 files of 800 MiB each (totaling 100 GiB). We then use Sysbench to issue continuous write requests with request sizes of 4 KiB, 8 KiB, 16 KiB, 32 KiB, and 64 KiB. For each request size, we run the workload for 10 minutes. Unless otherwise specified, we use one thread to issue writes in asynchronous mode with direct I/O, and each run ends with an fsync to ensure data persistence. Before each run, we clear and reformat the block device as an EXT4 file system. This workload stresses small, frequent writes and highlights both the throughput and metadata efficiency of different schemes.
Key-value store workloads. We evaluate ZapRAID when it hosts RocksDB (v9.7.2) [58] and use its benchmarking tool db_bench [53] to issue key-value operations. We focus on two settings of the fillrandom workload: (i) the default fillrandom, which inserts randomly ordered keys into an empty database and triggers background compaction, and (ii) fillrandom with compaction disabled (bulk loading), which sequentially writes data without compaction overhead. We configure the database with approximately 1.64 million key-value pairs, each with a value size of 64 KiB, and disable compression. We use eight threads per client and run a total of eight clients concurrently, resulting in 64 writer threads in total. The workload writes about 100 GiB of data in total, evenly distributed across all clients. For each run, we reset the volume and then issue one of the two fillrandom workloads. These two workload settings capture both the compaction overhead of LSM-tree maintenance and the pure bulk-loading behavior without compaction.
Concurrent workloads. We combine two types of write-intensive workloads with different write patterns to evaluate ZapRAID. Specifically, we use Sysbench to generate small file writes, while simultaneously hosting MySQL and running the oltp_write_only benchmark, which issues large writes with fewer than 10% reads. This setting emulates a virtualized environment in which multiple applications coexist in an SSD RAID array [14]. For the small file writes, we use Sysbench to continuously issue 4 KiB write requests. For the oltp_write_only workload, we use Sysbench to populate the MySQL database with 8 tables and 10 million rows, and then execute DELETE, INSERT, and UPDATE queries using 64 threads. Before each run, we clear and reformat the block device as an EXT4 file system, and then execute both benchmarks simultaneously for 10 minutes.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) File writes, single (4 KiB) (b) File writes, single (16 KiB) (c) File writes, multiple $(N_{s},N_{l})=(2,2)$ (d) File writes, multiple $(N_{s},N_{l})=(1,3)$
Refer to caption Refer to caption Refer to caption Refer to caption
(e) Key-value (with compaction), single (f) Key-value (with compaction), multiple (g) Key-value (bulk load, no compaction), single (h) Key-value (bulk load, no compaction), multiple
Refer to caption Refer to caption Refer to caption Refer to caption
(i) Concurrent: file writes (4 KiB), single (j) Concurrent: file writes (4 KiB), multiple (k) Concurrent: oltp_ write_only, single (l) Concurrent: oltp_ write_only, multiple
Figure 6: Exp#1 (Write performance): file writes (figures (a)-(d)), key-value store workloads (figures (e)-(h)), and concurrent workloads (figures (i)-(l)).
Exp#1 (Write performance). We first evaluate the write performance of ZapRAID and compare it against ZoneWrite-Only and ZoneAppend-Only under different open-segment settings.
File writes. Figures 6(a)-6(d) present the throughput results of sequential file writes under different open-segment settings. For a single open segment with 4 KiB chunks (Figure 6(a)), ZapRAID achieves throughput comparable to ZoneAppend, while outperforming ZoneWrite-Only by 38.6%. At larger request sizes, the performance gap between ZapRAID and ZoneWrite-Only increases significantly, reaching 60.9%, 73.0%, 71.2%, and 76.8% at 8 KiB, 16 KiB, 32 KiB, and 64 KiB, respectively. The gains mainly stem from ZapRAID’s ability to exploit intra-zone parallelism under a small chunk size, which is consistent with our observation in Figure 2(a) in §2.2 for a single open segment with 4 KiB chunks.
At small request sizes (e.g., 4 KiB), ZapRAID’s improvement is relatively modest due to the system overhead from file system metadata updates, as well as kernel I/O stack and NVMe-oF’s TCP stack operations. As the request size becomes larger, such overhead becomes amortized. Since the chunk size is 4 KiB, large requests are still issued to small chunks. ZapRAID can deliver progressively larger throughput gains over ZoneWrite-Only via intra-zone parallelism.
Settings Request sizes Findings Reasons
Single segment, 4 KiB chunks 4-8 KiB Modest improvements over ZoneWrite-Only Reasons A and B
16-64 KiB Significant improvements over ZoneWrite-Only Reason A
Single segment, 16 KiB chunks 4-8 KiB Marginal difference Reason C
16-64 KiB
Multiple segments, $(N_{s},N_{l})=(2,2)$ 4-8 KiB Marginal difference Reason D
16-64 KiB Marginal difference Reason C
Multiple segments, $(N_{s},N_{l})=(1,3)$ 4-8 KiB Modest improvements over ZoneWrite-Only Reasons A and B
16-64 KiB Marginal difference Reason C
Reason A: Exploiting intra-zone parallelism with small chunks; see Figures 2(a) and 2(b) in §2.2.
Reason B: System overhead of file system metadata updates, I/O stacks, and NVMe-oF’s TCP stacks.
Reason C: Large writes have similar throughput in Zone Append and Zone Write; see Figure 2(c) in §2.2.
Reason D: Small writes are distributed across zones and have similar throughput in Zone Append and Zone Write; see Figures 2(a) and 2(b) in §2.2.
Table 1: Exp#1 (Write performance). Explanations of the throughput results for file writes across different request sizes and segment settings.
For a single open segment with 16 KiB chunks and multiple open segments with $(N_{s},N_{l})=(2,2)$ (Figures 6(b) and 6(c), respectively), the performance differences among the three schemes are small. For the former, the 4 KiB writes are absorbed into 16 KiB chunks and issued as large writes, which are handled comparably by Zone Append and Zone Write (Figure 2(c) in §2.2); for the latter, the 4 KiB writes are distributed across two zones, where Zone Append and Zone Write deliver similar performance (Figure 2(a) in §2.2). For multiple open segments with $(N_{s},N_{l})=(1,3)$ (Figure 6(d)), ZapRAID shows a modest advantage at small request sizes (subject to the system overhead), outperforming ZoneWrite-Only by 20.2% and 16.6% at 4 KiB and 8 KiB, respectively, while the difference narrows at large request sizes since large writes yield comparable throughput across schemes.
Table 1 summarizes the reasons above to provide a consolidated view across different request sizes and open segment settings.
Note that across all settings (Figures 6(a)-6(d)), small requests of 4 KiB and 8 KiB can only achieve the throughput of at most 1,061.50 MiB/s and 1,789.36 MiB/s, respectively. The throughput is constrained by the overhead of file system metadata updates, kernel I/O stack operations, and TCP stack operations in NVMe-oF. In Exp#1, each request must traverse both the file system and TCP/IP stacks, which incur substantial processing overhead. In contrast, Exp#5 (§5.4) issues requests directly to the SPDK target without traversing the TCP stack, and the throughput can reach up to 1,599 MiB/s and 3,012 MiB/s under 4 KiB and 8 KiB writes, respectively. The TCP overhead in NVMe-oF, even with TCP loopback, has also been reported by the SPDK community [67]. Such overhead becomes less significant for large request sizes.
Key-value store workloads. Figures 6(e)-6(h) show the throughput results of RocksDB using db_bench under two configurations: the fillrandom workload with compaction enabled and the bulk-loading workload without compaction.
Under fillrandom with compaction (Figures 6(e) and 6(f)), ZapRAID outperforms ZoneWrite-Only by 41.7% under a single open segment with 4 KiB chunks, while the three schemes perform nearly the same as under a single open segment with 16 KiB chunks (Figure 6(e)). For multiple open segments with $(N_{s},N_{l})=(2,2)$ and $(N_{s},N_{l})=(1,3)$, the throughput difference across the three schemes remains small (Figure 6(f)).
For bulk loading without compaction (Figures 6(g) and 6(h)), ZapRAID shows larger benefits at small request sizes. It achieves 51.7% higher throughput than ZoneWrite-Only under a single open segment with 4 KiB chunks, while the advantage narrows to 5.7% under a single open segment with 16 KiB chunks. As with the compaction-enabled workload, the multiple open segments with $(N_{s},N_{l})=(2,2)$ and $(N_{s},N_{l})=(1,3)$ yield comparable throughput across all schemes. Since compaction is disabled during bulk loading, read overheads are eliminated, making the performance improvement of ZapRAID more pronounced.
We observe that the performance gaps across the three schemes are small for the configurations with a large chunk size (e.g., 16 KiB) or sufficient open segments (e.g., four). RocksDB exhibits both small writes (e.g., write-ahead log (WAL) writes) and large writes (e.g., flushes and compactions), while large writes dominate. Both Zone Write and Zone Append primitives are expected to have close performance in such configurations (see Figure 2 in §2.2).
Concurrent workloads. Figures 6(i) and 6(j) show the write throughput results for small file writes in concurrent workloads. ZapRAID achieves 26.3% and 34.5% higher throughput than ZoneWrite-Only under a single open segment with 4 KiB chunks (Figure 6(i)) and four open segments with $(N_{s},N_{l})=(1,3)$ (Figure 6(j)), respectively. In both settings, all small writes are directed to the single small-chunk segment. ZapRAID outperforms ZoneWrite-Only by issuing Zone Append. On the other hand, for a single open segment with 16 KiB chunks and four open segments with $(N_{s},N_{l})=(2,2)$, all three schemes have similar performance, due to the same reasons as explained in the file writes workloads (see above).
Figures 6(k) and 6(l) depict the throughput of the oltp_write_only benchmark in transactions per second (TPS). For a single open segment with 4 KiB chunks (Figure 6(k)), ZapRAID achieves 32.9% higher throughput than ZoneWrite-Only and has similar throughput to ZoneAppend-Only. For other open-segment settings, all three schemes show similar performance. In oltp_write_only, where large writes dominate, we expect similar performance across all three schemes for the configurations with a large chunk size and sufficient open segments (see the explanations above for key-value store workloads).
Exp#2 (Degraded reads). We evaluate degraded read performance by first loading data into the system, failing one disk to emulate degraded mode, and finally issuing read requests based on the file writes and key-value store workloads. We consider two workloads, derived from the file writes and key-value store workloads, respectively: (i) Sysbench with readrandom and (ii) db_bench with readrandom. For the former, we load a 100 GiB dataset using Sysbench, fail one device, and perform degraded random reads continuously for 10 minutes using Sysbench in readrandom mode. For the latter, we use db_bench to first load a 100 GiB database with 64 KiB values, fail one device, and then issue 20,000 degraded readrandom requests using db_bench.
[…]
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Sysbench, single (b) Sysbench, multiple (c) db_bench, single (d) db_bench, multiple
Figure 7: Exp#2 (Degraded reads): Sysbench (figures (a) and (b)) and db_bench (figures (c) and (d)).
[…]
File writes Single-4 KiB Single-16 KiB $(N_{s},N_{l})=(2,2)$ $(N_{s},N_{l})=(1,3)$
(ZoneAppend-Only / ZapRAID)
L2P Table 100.00 / 100.00 100.00 / 100.00 100.00 / 100.00 100.00 / 100.00
Segment table 0.0008 / 0.0008 0.0008 / 0.0008 0.0008 / 0.0008 0.0008 / 0.0008
Compact stripe table 106.68 / 35.56 26.67 / 8.89 26.67 / 0.00 26.67 / 0.00
Total 206.68 / 135.56 126.67 / 108.89 126.67 / 100.00 126.67 / 100.00
Key-value store Single-4 KiB Single-16 KiB $(N_{s},N_{l})=(2,2)$ $(N_{s},N_{l})=(1,3)$
(ZoneAppend-Only / ZapRAID)
L2P table 77.55 / 77.55 77.55 / 77.55 77.55 / 77.55 77.55 / 77.55
Segment table 0.0006 / 0.0006 0.0006 / 0.0006 0.0006 / 0.0006 0.0006 / 0.0006
Compact stripe table 81.58 / 27.19 20.40 / 6.80 20.40 / 0.00 20.40 / 0.00
Total 159.13 / 104.74 97.95 / 84.35 97.95 / 77.55 97.95 / 77.55
Table 2: Exp#3 (Memory usage for index management). Breakdown of memory usage (MiB) for index management in ZoneAppend-Only and ZapRAID under the file writes and key-value store workloads.
Exp#3 (Memory usage for index management). To evaluate memory usage for index management, we consider the file writes and key-value store workloads. We assume that the L2P table is entirely in memory, without being offloaded (§3.1). Table 2 reports the breakdown of memory usage for index management (including the L2P table, segment table, and compact stripe table) and the total memory usage.
For file writes, we run Sysbench with a request size of 64 KiB to generate a 100 GiB dataset, while for key-value store workloads, we run db_bench with compaction enabled to write a total of 100 GiB of data across eight clients with 64 KiB requests. We compare ZoneAppend-Only and ZapRAID, since they both maintain stripe-to-chunk mappings under Zone Append, while ZoneWrite-Only does not issue Zone Append and hence has no compact stripe table memory consumption. The workload settings are identical to those used in Exp#1.
For file writes, the L2P table and segment table sizes stay constant across configurations (100.00 MiB and 0.0008 MiB, respectively), as they depend only on the total volume of data written. The differences come from the compact stripe table. For a single open segment with 4 KiB chunks, the compact stripe table decreases from 106.68 MiB (ZoneAppend-Only) to 35.56 MiB (ZapRAID) (i.e., 66.7% reduction), so ZapRAID reduces the total memory usage of ZoneAppend-Only from 206.68 MiB to 135.56 MiB (34.4%). We emphasize that ZapRAID’s memory reduction (66.7%) of the compact stripe table compared to ZoneAppend-Only is critical, since it significantly reduces the degraded read overhead (Exp#2).
ZapRAID also shows memory usage reduction in other settings. For a single open segment with 16 KiB chunks, the compact stripe table drops from 26.67 MiB to 8.89 MiB (66.7%), and ZapRAID reduces the total memory usage of ZoneAppend-Only from 126.67 MiB to 108.89 MiB (14.0%). For multiple open segments with $(N_{s},N_{l})=(2,2)$ and $(N_{s},N_{l})=(1,3)$, ZapRAID directs all 64 KiB writes to large segments, so it reverts to ZoneWrite and does not require a compact stripe table. Thus, ZapRAID reduces the compact stripe table size from 26.67 MiB to 0.00 MiB, and the total memory usage is reduced from 126.67 MiB to 100.00 MiB (21.1%), compared to ZoneAppend-Only.
For key-value store workloads, the trends are similar. The L2P and segment table sizes remain fixed at 77.55 MiB and 0.0006 MiB, respectively, while the compact stripe table and total memory usage show clear savings in ZapRAID. For example, for a single open segment with 4 KiB chunks, ZapRAID reduces the compact stripe table size from 81.58 MiB to 27.19 MiB (66.7%), reducing the total memory usage from 159.13 MiB to 104.74 MiB (34.2%), compared to ZoneAppend-Only.
Summary of real-application experiments. Compared to ZoneWrite-Only, ZapRAID excels in small chunk sizes and increases the write throughput by up to 76.8% (in 64 KiB writes under Sysbench). A trade-off is that ZapRAID introduces a compact stripe table, but it limits the memory usage of the compact stripe table (as well as the degraded read overhead) through the group-based data layout.
Compared to ZoneAppend-Only, ZapRAID achieves nearly identical write throughput (Exp#1), while significantly improving degraded read performance (Exp#2) and reducing memory usage (Exp#3). For example, in Sysbench file writes under a single open segment with 4 KiB chunks, ZapRAID’s write throughput is within 2.42% of ZoneAppend-Only’s, while reducing the compact stripe table size by 66.7% and the total memory usage for index management by 34.4%. Similarly, in the db_bench fillrandom workload under a single segment with 4 KiB chunks, ZapRAID’s write throughput is within 2.20% of ZoneAppend-Only’s, while reducing the compact stripe table size by 66.7% and the total memory usage for index management by 34.2%. By limiting the compact stripe table size, ZapRAID also significantly outperforms ZoneAppend-Only in degraded reads.
5.3 Trace-Driven Experiments (RQ2)
Exp#4 (Cloud block storage workloads). The original Alibaba Cloud traces cover 1,000 volumes [41]. We identify that 684 out of the 1,000 volumes have more than 60% of small write requests of no more than 4 KiB and more than 25% of large write requests of at least 16 KiB (§2.2). Among the 684 volumes, we choose 20 of them whose amounts of write traffic range between 200 GiB and 240 GiB. Note that the selected 20 volumes are dominated by writes: in 18 out of 20 volumes, more than 88.3% of requests are writes, while in all volumes, more than 67.0% of requests are writes. We conduct our evaluation over five runs; in each run, we evaluate the average throughput and 95th-percentile latency of all read and write operations when running each scheme over the 20 selected volumes.
Figures 8(a) and 8(b) show the throughput results. For a single open segment with 4 KiB and 16 KiB chunks (Figure 8(a)), the throughput of ZapRAID is 69.4% and 6.4% higher than that of ZoneWrite-Only, respectively, and is close to that of ZoneAppend-Only; For multiple open segments $(N_{s},N_{l})=(2,2)$, all three schemes have similar throughput, while for $(N_{s},N_{l})=(1,3)$, ZapRAID has 25.3% higher throughput than ZoneWrite-Only and similar throughput to ZoneAppend-Only.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Throughput (single segment) (b) Throughput (multiple segments) (c) p95 latency (single segment) (d) p95 latency (multiple segments)
Figure 8: Exp#4 (Cloud block storage workloads). Average throughput and 95 latency results of all 20 volumes in Alibaba Cloud traces.
Figures 8(c) and 8(d) show the 95th-percentile latencies. For a single open segment with 4 KiB and 16 KiB chunks (Figure 8(c)), ZapRAID has 56.6% lower and 39.6% higher 95th-percentile latencies than ZoneWrite-Only under the chunk sizes of 4 KiB and 16 KiB, respectively, and comparable 95th-percentile latencies with ZoneAppend-Only. For multiple open segments (Figure 8(d)), all three schemes have similar latencies for $(N_{s},N_{l})=(2,2)$, while ZoneWrite-Only has a higher 95th-percentile latency for $(N_{s},N_{l})=(1,3)$ as its performance is bottlenecked by the small writes that are forwarded to the only small-chunk open segment.
Volume No. 1 2 3 4 5 6 7 8 9 10
Large writes 1.63 1.64 1.67 1.70 1.75 1.76 1.77 1.78 1.83 1.86
Thpt (MiB/s) $($$N_{s},N_{l}$)=(1,3)) ZoneWrite-Only 2,486.4 2,341.2 2,323.1 2,401.3 2,403.7 2,333.7 2,311.0 2,360.3 2,344.8 2,410.5
Volume No. 11 12 13 14 15 16 17 18 19 20
Large writes 1.92 1.95 3.42 3.70 4.49 10.33 16.83 20.95 21.11 24.89
Thpt (MiB/s) $($$N_{s},N_{l}$)=(1,3)) ZoneWrite-Only 2,372.7 2,399.1 2,412.2 2,439.7 2,593.6 2,637.1 2,703.2 2,773.4 2,862.1 2,825.4
Table 3: Exp#4 (Cloud block storage workloads). Statistics of the 20 individual volumes in Alibaba Cloud traces.
We also show the statistics of individual volumes by presenting the ratios of numbers of small writes (i.e., 4 KiB) and large writes (i.e., at least 16 KiB) over the total number of write requests as well as the throughput of each volume. For brevity, since all three schemes have close throughput under one single open segment with 16 KiB chunks and four open segments with $(N_{s},N_{l})=(2,2)$, we only present the results of one single open segment with 4 KiB chunks and four open segments with $(N_{s},N_{l})=(1,3)$. Also, we only focus on the results under ZoneWrite-Only and ZapRAID to show their variations across the volumes.
Table 3 shows the results. We sort the volumes by the ratio of large write requests. For one single open segments under 4 KiB chunks, the throughput is close over all volumes for each scheme. For four open segments with $(N_{s},N_{l})=(1,3)$, ZoneWrite-Only shows lower throughput for a lower ratio of large writes; for example, the throughput results of Volumes 2 and 20 are 2,341.2 MiB/s and 2,825.4 MiB/s with large-write ratios of 1.64% and 24.89%, respectively. The reason is that with fewer large writes, more write requests are written to the only $N_{s}=1$ small-chunk open segment, and ZoneWrite-Only fails to utilize intra-zone parallelism to issue writes efficiently. In contrast, ZapRAID maintains higher throughput with an increase of 14.7-40.8% compared with ZoneWrite-Only under $(N_{s},N_{l})=(1,3)$, across all 20 volumes.
5.4 Microbenchmarks (RQ3)
@@
We examine the performance of ZapRAID in writes, normal reads, degraded reads, crash recovery, and full-drive recovery. In our default setup, we configure the logical space capacity as 200 GiB and set $G=256$ for ZapRAID in its group-based data layout. First, we focus on one open segment with various chunk sizes in Exp#5-Exp#10. We do not consider garbage collection (by limiting the total write size to be less than the logical space capacity) and keep the L2P table entirely in memory. We also consider the impact of various group sizes (Exp#7), the impact of RAID schemes (Exp#8), as well as the crash recovery and full-drive recovery performance with different logical space capacities (Exp#9). Furthermore, we consider multiple open segments in Exp#11-Exp#13, including the write performance under multiple open segments (Exp#11), the overhead of garbage collection (Exp#12), and the overhead of L2P table offloading (Exp#13).
@@
Exp#5 (Write performance). We first evaluate the write performance of ZapRAID on a single open segment. We use Flexible IO Tester (FIO) (v3.35) [17] to generate a write-only workload that issues random writes of 64 GiB of data. We set the queue depth as 64 to saturate the system parallelism (Exp#10). We compare ZapRAID with RAIZN-SPDK, ZoneWrite-Only, and ZoneAppend-Only.
@@
First, we issue write requests of 4 KiB, 8 KiB, and 16 KiB, where the request size is the same as the chunk size. Figures 9(a)-9(c) show the throughput as well as the median and 95th-percentile latencies. Compared with ZoneWrite-Only, for 4 KiB and 8 KiB writes, ZapRAID achieves 72.8% and 77.2% higher write throughput, 43.8% and 44.5% lower median latency, and 36.4% and 34.1% lower 95th-percentile latency, respectively. This shows that ZapRAID improves the write performance by exploiting the intra-zone parallelism using Zone Append for small writes. On the other hand, for 16 KiB writes, ZapRAID has similar write throughput and median latency, but its 95th-percentile latency is $3.80\times$ compared with that of ZoneWrite-Only. A possible reason is that Zone Append has high computational overhead in the ZNS SSD firmware, which incurs higher delays in processing individual Zone Append commands.
[…]
@@
Compared with ZoneAppend-Only, ZapRAID achieves similar throughput and latencies for all write sizes, even though ZapRAID needs to wait for the completion of all Zone Append commands in a stripe group before issuing the writes for the next stripe group. Note that ZapRAID has much less memory usage (§3.2) and much lower degraded read latency (Exp#7) than ZoneAppend-Only.
@@
Compared with RAIZN-SPDK, all other three schemes show much higher performance on a single segment. The reasons are two-fold. First, RAIZN does not support Zone Append for intra-zone parallelism. Second, it uses partial parity updates to acknowledge individual write requests instead of issuing full-stripe writes as in the other three schemes. This limits the concurrency of write requests and hence the overall write performance. In Exp#11, we further compare ZapRAID with RAIZN-SPDK under multiple open segments.
Figure 9(d) shows the memory usage for stripe management per GiB of written data. Here, we focus on ZoneAppend-Only and ZapRAID, as they both need to maintain mappings between stripes and chunk locations under Zone Append (note that ZoneWrite-Only has no compact stripe table as it does not issue Zone Append). For 4 KiB writes, ZoneAppend-Only consumes 1,018.2 KiB of memory, while ZapRAID consumes 339.4 KiB only (i.e., 66.7% reduction) due to its compact stripe table design. For a (3+1)-RAID-5 array in which each drive has $Z=$ 3,690 zones and the zone capacity is 1,077 MiB, the maximum memory size for ZoneAppend-Only is $\frac{1018.2\times 3\times 3690\times 1077}{1024\times 1024\times 1024}=11.3$ GiB, while ZapRAID consumes 3.77 GiB only (§3.2). The result suggests that ZapRAID greatly reduces the memory footprints for stripe management compared with ZoneAppend-Only. Note that larger chunks imply a smaller compact stripe table since each segment has fewer chunks.
Refer to caption Refer to caption
(a) Write throughput (b) Median write latency
Refer to caption Refer to caption
(c) p95 write latency (d) Memory usage (per GiB of data) for compact stripe table
@@
Figure 9: Exp#5 (Write performance on a single open segment). We show the results when the request size is equal to the chunk size. For figure (d), we show ZoneAppend-Only and ZapRAID only.
Refer to caption Refer to caption
(a) Write throughput (b) Median write latency
Refer to caption Refer to caption
(c) p95 write latency (d) Memory usage (per GiB of data) for compact stripe table
Figure 10: Exp#5 (Write performance on a single open segment). We show the results when the request size is fixed as 4 KiB.
@@
We also issue write requests of a fixed size of 4 KiB, while varying the chunk size, to show the impact of the chunk size on write performance. Figure 10 shows the results. ZapRAID still maintains the high throughput and low median latencies for various chunk sizes, but its 95th-percentile latency for 16 KiB chunks is 50.7% higher than that for 8 KiB chunks. Similar results also appear in ZoneAppend-Only. The result indicates that Zone Append is mainly beneficial for small writes. It also justifies why we choose Zone Write over Zone Append in hybrid data management (§3.3).
@@
Exp#6 (Normal and degraded read performance). We compare the performance of both normal reads and degraded reads in Log-RAID and ZapRAID. We vary the chunk size and set the read size to be the same as the chunk size. For each chunk size, we first fill up the 200 GiB logical address space to ensure that every logical block being read is indeed physically stored. We then use FIO to generate a read-only workload that issues random reads to 16 GiB of data using the randread option in FIO. We set the queue depth as one, so as to focus on the performance of individual read requests and exclude the interference among requests. We consider three cases: normal reads (NR), degraded reads under the static mapping in Log-RAID (DR-Log-RAID) (§2.3), and degraded reads under the group-based data layout in ZapRAID (DR-ZapRAID). Note that both Log-RAID and ZapRAID have the same workflow for normal reads. To evaluate degraded reads, we fail a drive and issue reads to the lost blocks of the failed drive.